### v1

@@ 3 Main Results
Focusing on the most informative equilibrium in constant threshold strategies, $I(\theta;\mathbf{h}_{\delta})$ is a continuous and strictly increasing function of $\delta$.
Proof.
For any $n\geq 2$ fix $\delta^{\prime}<\delta$ so that $\delta^{\prime},\delta\in(\delta_{n+1},\delta_{n}]$. We first show that $I(\theta;\mathbf{h}_{\delta^{\prime}})\leq I(\theta;\mathbf{h}_{\delta})$. Observe that both $\mathbf{h}_{\delta}$ and $\mathbf{h}_{\delta^{\prime}}$ involve disclosing signals $\{s_{1},...,s_{n}\}$. We refer to the sequence of signals after the last communication failure (including the ones concealed) as the complete history of game. Since the expected length of the complete history in the steady state equals $\delta/(1-\delta)$, the expected number of signals disclosed is increasing in $\delta$. So if $\delta^{\prime}<\delta$ the random variable $\mathbf{h}_{\delta^{\prime}}$ can be represented as a garbling of the random variable $\mathbf{h}_{\delta}$ and hence $I(\theta;\mathbf{h}_{\delta^{\prime}})\leq I(\theta;\mathbf{h}_{\delta})$. Observe also that for such $\delta$, $I(\theta;\mathbf{h}_{\delta})$ varies continuously since $\delta/(1-\delta)$ is continuous in $\delta$.
Fix any cutoff $\delta_{n}$. By definition, signals $\{s_{1},...,s_{n}\}$ are disclosed in the most informative constant threshold equilibrium when $\delta$ is slightly below $\delta_{n}$ and signals $\{s_{1},...,s_{n-1}\}$ are disclosed in the most informative constant threshold equilibrium when $\delta$ is slightly above $\delta_{n}$ At the cutoff $\delta_{n}$, the incentive constraint to disclose signal $n$ is binding, so that by equation (3.8) we have
But because of the binding constraint, disclosing an additional realization of $s_{n}$ does not change the likelihood ratio of the public belief. Hence, when $\delta=\delta_{n}$, we have $\pi(h_{1},...,h_{n-1},0,...,0)=\pi(h_{1},...,h_{n-1},h_{n},0,...,0)$ for any $h_{n}\in\mathbb{N}$ and $(h_{1},...,h_{n-1})\in\mathbb{N}^{n-1}$. Therefore, we have $I(\theta;\mathbf{h}^{(n)}_{\delta_{n}})=I(\theta;\mathbf{h}^{(n-1)}_{\delta_{n}})=I(\theta;\mathbf{h}_{\delta_{n}})$ and $I(\theta;\mathbf{h}_{\delta})$ is continuous at $\delta=\delta_{n}$. ∎
We conclude this section by discussing our refinements and comparing them with the ones in Guttman, Kremer, and Skrzypacz (2014), Hart, Kremer, and Perry (2017), and Rappoport (2025). First, Guttman, Kremer, and Skrzypacz (2014) define monotone equilibria and threshold equilibria when a sender can receive up to two signals in two periods and can decide what to disclose and when. We view our notions of monotone beliefs and threshold strategies as natural extensions of their refinements to our setting where the sender can have any number of signals.
Second, our threshold strategy refinement and our monotone belief refinement do not imply one another. For example, strategies where agents disclose only one realization of a signal even when multiple of them are available are not threshold strategies but are allowed in equilibria with monotone belief maps. Agents’ incentives to use a threshold strategy $J:H\times S\rightarrow H$ imply that they have no incentive to conceal any signal with an index below $J(\mathbf{h},s)$ at any $(\mathbf{h},s)$. Deviations where they conceal some of those signals will also reduce the length of the next agent’s history (conditional on communication succeeds). The monotonicity refinement on the belief map has no bite in ruling out such deviations as it only imposes restrictions on histories with the same length.
@@
We explain why there is no strict equilibrium that satisfies both our threshold strategy refinement and their (P0) requirement. The conflict between our monotone belief refinement and their (P0) requirement follows from the same logic. From Corollary 1, we know that if $\delta$ is large, then in every strict equilibrium in threshold strategies, agents will disclose $\mathbf{0}$ when their history is $\mathbf{0}$ and their private signal is $s_{2}$. According to Bayes rule, their belief about the state upon observing $\mathbf{0}$ is no more than their prior $\pi_{0}$. If such an equilibrium also satisfies (P0), then given that $\mathbf{1}_{2}$ occurring with zero probability, $\pi(\mathbf{1}_{2})$ must equal the agent’s posterior belief when they know that their predecessor’s history is $\mathbf{0}$ and their private signal is $s_{2}$. Since $L_{2}>1$, this leads to a posterior belief that is strictly greater than the posterior belief from disclosing nothing, which contradicts our earlier conclusion that agents will disclose $\mathbf{0}$ upon receiving private signal $s_{2}$ at history $\mathbf{0}$.
Fourth, our refinements are neither stronger nor weaker relative to the receiver’s optimal equilibrium, which is the refinement used in Rappoport (2025). First, the equilibrium where the sender never discloses any signal satisfies both of our refinements. However, it is not the receiver-optimal equilibrium as the one where the sender only discloses $s_{1}$ leads to higher receiver welfare. Second, suppose $L_{1}>L_{2}>1>$... but $\overline{f_{1}}$ and $\underline{f_{1}}$ are much smaller relative to $\overline{f_{2}}$ and $\underline{f_{2}}$. When $\delta$ is close to $1$, only signal $s_{1}$ can be disclosed in any strict equilibrium that satisfies any of our two refinements, whereas the receiver is strict better off in the equilibrium where only signal $s_{2}$ is disclosed. The latter equilibrium exists for any $\delta$ according to our Proposition 1.
@@ 4 Proof of Theorem 1
We begin by stating several useful properties of strict equilibria of our game. We start from a gradualism property.
@@
Second, when an agent decides what information to disclose, his objective is to maximize his immediate successor’s belief. Since he has the option to ignore his private signal and reveal his history $\mathbf{h}$, from which he can induce the same public belief when communication succeeds, the history he discloses in equilibrium must lead to a weakly higher belief compared to $\pi(\mathbf{h})$. This leads to an improvement principle similar to the one in Banerjee and Fudenberg (2004).
@@
For any equilibrium and any two adjacent periods $t$ and $t+1$, if the realized history in period $t$ is $\mathbf{h}^{t}$ and the history disclosed by period $t$ agent is $\mathbf{h}^{t+1}$, then $\pi(\mathbf{h}^{t+1})\geq\pi(\mathbf{h}^{t})$. If this equilibrium is strict, then $\pi(\mathbf{h}^{t+1})>\pi(\mathbf{h}^{t})$ as long as $\mathbf{h}^{t+1}\neq\mathbf{h}^{t}$.
@@
For any strict equilibrium, any history $\mathbf{h}$ that occurs with positive probability in this equilibrium, and any $\mathbf{h^{\prime}}<\mathbf{h}$, history $\mathbf{h^{\prime}}$ will not occur with positive probability after history $\mathbf{h}$ but before the next communication failure.
@@
Using these properties, one can see why Corollary 1 is a direct implication of Theorem 1. Since only $s_{1}$ can be disclosed when $\delta$ is large enough and any strict equilibrium needs to respect the no-turning-back property, there are only two threshold strategies that meet both requirements: one where no signal is disclosed and one where agents disclose all realized $s_{1}$. [^8] The no-turning-back property rules out strategies where at histories with a positive number of $s_{1}$, agents set the disclosure threshold to be 0, i.e., conceal all realized signals. Strict equilibria with monotone belief maps turn out to be more permissive: any strategy where agents disclose up to $n$ realizations of $s_{1}$ can be part of such equilibria as long as the belief map $\pi$ assigns a low value at any history with length more than $n$.
@@
Let $H(j)\subset H_{\sigma}$ denote the set of positive-probability histories under $\sigma$ such that the $j$th entry is $1$. Let $H(j-1)\subset H(j)$ denote the subset of $H(j)$ such that a history $\mathbf{h}$ belongs to H(j1)H(j-1) if and only if it minimizes $h_{1}+...+h_{j-1}$ among all vectors in $H(j)$. Iteratively define H(j2),,H(1)H(j-2),...,H(1) with $H(1)\subset H(2)\subset...\subset H(j-2)\subset H(j-1)$ such that for every $i\in\{1,2,...,j-2\}$, $\mathbf{h}$ belongs to $H(i)$ if and only if it minimizes $h_{1}+...+h_{i}$ among all vectors in $H(i+1)$. This leads to a unique vector in $H(1)$, which we denote by $\mathbf{h^{*}}\equiv(h_{1}^{*},...,h_{l}^{*})$.
Equilibria in Threshold Strategies:
Suppose for some $\delta$ is close enough to $1$, there exists a strict equilibrium where agents use threshold strategy $\sigma$ under which $j\geq 2$. We establish the following lemma, which shows that $\mathbf{h^{*}}$ can only originate from $\mathbf{h^{*}}-\mathbf{1}_{j}$.
@@
Suppose in some strict equilibrium, agents use a threshold strategy. If an agent’s history is $\mathbf{h^{*}}$, then his immediate predecessor’s history is either $\mathbf{h^{*}}$ or $\mathbf{h^{*}}-\mathbf{1}_{j}$.
The proof is in Appendix B.
$\mu_{\theta}(\mathbf{h^{*}})=\mu_{\theta}(\mathbf{h^{*}}-\mathbf{1}_{j})\cdot\delta f_{\theta,j}\cdot\sum_{k=0}^{\infty}\delta^{k}(1-\sum_{i=1}^{j}f_{\theta,i})^{k}$
$\mu_{\theta}(\mathbf{h^{*}}-\mathbf{1}_{j})\cdot\frac{\delta f_{\theta,j}}{1-\delta(1-\sum_{i=1}^{j}f_{\theta,i})}.$ (4.1)
After receiving private signal $s_{j}$ at history $\mathbf{h^{*}}-\mathbf{1}_{j}$, an agent has an incentive to disclose history $\mathbf{h^{*}}$ to his successor only if $\pi(\mathbf{h^{*}})\geq\pi(\mathbf{h^{*}}-\mathbf{1}_{j})$, which according to (4.1), translates into
When $\delta$ is close to $1$, the above inequality fails for any $j\geq 2$, which leads to a contradiction.
Equilibria with Monotone Belief Maps:
Suppose for some $\delta$ is close enough to $1$, there exists a strict equilibrium with a monotone belief map $\pi$ under which $j\geq 2$. The monotonicity of $\pi$ implies that $\pi(\mathbf{h^{*}})\leq\pi(\mathbf{h})$ for any $\mathbf{h}\in H(j-1)$. We establish the following lemma.
  Lemma 2.
Suppose the belief map is monotone in some strict equilibrium. If an agent’s history is $\mathbf{h^{*}}$, then his immediate predecessor’s history is either $\mathbf{h^{*}}$ or $\mathbf{h^{*}}-\mathbf{1}_{j}$.
The proof is in Appendix C.
For every $\mathbf{h}\in H_{\sigma}$, let $D(\mathbf{h})\subset\{1,2,...,l\}$ be such that $i\in D(\mathbf{h})$ if and only if the disclosed vector after receiving private signal $s_{i}$ at history $\mathbf{h}$ is not weakly dominated by $\mathbf{h}$, or equivalently, the number of signal $s_{i}$ disclosed is more than the $i$th entry of $\mathbf{h}$. Let $\overline{d}(\mathbf{h})\equiv\max D(\mathbf{h})$.
Since the equilibrium is strict, we know that $\pi(\mathbf{h})>\pi(\mathbf{h^{\prime}})$ for every $\mathbf{h}\in H_{\sigma}$ and $\mathbf{h^{\prime}}<\mathbf{h}$. Hence, if the agent receives private signal $s_{i}$ at history $\mathbf{h}$ with $i\notin D(\mathbf{h})$, he will disclose history $\mathbf{h}$ to his immediate successor, so the history will remain the same as long as communication succeeds. This together with Lemma 2 leads to

### v2

@@ 3 Main Results
The function $I(\theta;\mathbf{h}_{\delta})$ is continuous and strictly increasing in $\delta$.
The proof is in Appendix A.3.
We conclude this section by discussing our refinements and comparing them with the ones in the existing literature. First, Guttman, Kremer, and Skrzypacz (2014) define monotone equilibria and threshold equilibria when a sender can receive up to two signals in two periods. We view our notions of monotone equilibria and threshold equilibria as natural extensions of their refinements to our setting where the sender can have any number of signals.
Second, Corollary 1 implies that when $\delta$ is above some threshold $\delta^{*}$, for every strict threshold equilibrium, there exists a strict monotone equilibrium where agents use the same strategy but not vice versa. However, under general values of $\delta$, it is unclear whether our monotone belief refinement is more permissive than our threshold strategy refinement. Intuitively, agents’ incentives to use a threshold strategy $J:H\times S\rightarrow H$ imply that they have no incentive to conceal any signal with an index below $J(\mathbf{h},s)$ at any $(\mathbf{h},s)$. Deviations where they conceal some of those signals will also reduce the length of the next agent’s history (conditional on communication succeeds). The monotonicity refinement on the belief map has no bite in ruling out such deviations as it only imposes restrictions on histories with the same length. Section 5.1 will revisit the differences between our two refinements when we discuss the differences between our overlapping generations game and one-shot disclosure games between a sender and a receiver where the sender has an unknown number of signals.
@@
We explain why there is no strict equilibrium that satisfies both our threshold strategy refinement and their (P0) requirement when $\delta$ is large. The conflict between our monotone belief refinement and their (P0) requirement follows from the same logic. From Corollary 1, we know that in every strict threshold equilibrium, agents will disclose $\mathbf{0}$ when their history is $\mathbf{0}$ and their private signal is $s_{2}$. According to Bayes rule, the probability with which their belief assigns to $\overline{\theta}$ upon observing history $\mathbf{0}$ is no more than their prior belief $\pi_{0}$. If such an equilibrium also satisfies (P0), then given that history $\mathbf{1}_{2}$ occurring with zero probability, $\pi(\mathbf{1}_{2})$ must equal the agent’s posterior belief when they know that their predecessor’s history is $\mathbf{0}$ and their private signal is $s_{2}$. Since $L_{2}>1$, we have $\pi(\mathbf{1}_{2})>\pi_{0}>\pi(\mathbf{0})$. This contradicts our earlier conclusion that agents will disclose $\mathbf{0}$ upon receiving private signal $s_{2}$ at history $\mathbf{0}$.
Fourth, our refinements are neither stronger nor weaker relative to the receiver’s optimal equilibrium used in Rappoport (2025) and Gao (2025). First, the equilibrium where the sender never discloses any signal satisfies both of our refinements. However, it is not the receiver-optimal equilibrium as the one where the sender only discloses $s_{1}$ leads to higher receiver welfare. Second, suppose $L_{1}>L_{2}>1>$... but $\overline{f_{1}}$ and $\underline{f_{1}}$ are much smaller relative to $\overline{f_{2}}$ and $\underline{f_{2}}$. When $\delta$ is close to $1$, only signal $s_{1}$ can be disclosed in any strict equilibrium that satisfies any of our two refinements, whereas the receiver is strict better off in the equilibrium where only signal $s_{2}$ is disclosed. The latter equilibrium exists for any $\delta$ according to our Proposition 1.
@@ 4 Proof of Theorem 1
We state several useful properties of strict equilibria, without imposing the monotone belief refinement and the threshold strategy refinement. We start from a gradualism property.
@@
Second, when an agent decides what information to disclose, his objective is to maximize his immediate successor’s belief. Since he has the option to ignore his private signal and reveal his history $\mathbf{h}$, from which he will induce the same belief when communication succeeds, the history he discloses in equilibrium must lead to a weakly higher belief compared to $\pi(\mathbf{h})$. This leads to an improvement principle similar to the one in Banerjee and Fudenberg (2004).
@@
For any equilibrium and any $t\in\mathbb{Z}$, if the realized history in period $t$ is $\mathbf{h}^{t}$ and agent $t$ chooses to disclose $\mathbf{h}^{t+1}$ after observing some private signal at $\mathbf{h}^{t}$, then $\pi(\mathbf{h}^{t+1})\geq\pi(\mathbf{h}^{t})$. If the equilibrium is strict and $\mathbf{h}^{t+1}\neq\mathbf{h}^{t}$, then $\pi(\mathbf{h}^{t+1})>\pi(\mathbf{h}^{t})$. [^11] Our improvement principle is stated in terms of beliefs since for each agent in our model, maximizing his expected payoff from his successor’s action is equivalent to choosing $\mathbf{h}$ in order to maximize $\pi(\mathbf{h})$. The equivalence between maximizing payoffs and maximizing beliefs breaks down in one of our extensions, in which case a more general version of the improvement principle holds: Disclosing $\mathbf{h}^{t+1}$ leads to a weakly greater expected payoff for agent $t$ compared to disclosing $\mathbf{h}^{t}$, where the inequality is strict when the equilibrium is strict and $\mathbf{h}^{t+1}\neq\mathbf{h}^{t}$.
@@
In any strict equilibrium, for any history $\mathbf{h}$ that occurs with positive probability in this equilibrium as well as any $\mathbf{h^{\prime}}<\mathbf{h}$, history $\mathbf{h^{\prime}}$ will not occur after history $\mathbf{h}$ before the next communication failure.
@@
Using these properties, one can see why Corollary 1 is a direct implication of Theorem 1. Since only $s_{1}$ can be disclosed when $\delta$ is large enough and any strict equilibrium needs to respect the no-turning-back property, there are only two threshold strategies that meet both requirements: one where no signal is disclosed and one where agents disclose all realized $s_{1}$. [^12] The no-turning-back property rules out strategies where at histories that occur with positive probability and contain a positive number of $s_{1}$, agents set the disclosure threshold to be 0, i.e., conceal all realized signals. When $\delta$ is large enough, strict monotone equilibrium turns out to be a more permissive concept: any strategy where agents disclose up to $n$ realizations of $s_{1}$ can be part of such an equilibrium as long as the belief map $\pi$ assigns a low value at any history with length more than $n$.
@@
Let $H(j)\subset H_{\sigma}$ denote the set of positive-probability histories under $\sigma$ such that the $j$th entry is $1$. Let $H(j-1)\subset H(j)$ denote the subset of $H(j)$ such that a history $\mathbf{h}$ belongs to H(j1)H(j-1) if and only if it minimizes $h_{1}+...+h_{j-1}$ among all vectors in $H(j)$. Iteratively define H(j2),,H(1)H(j-2),...,H(1) with $H(1)\subset H(2)\subset...\subset H(j-2)\subset H(j-1)$ such that for every $i\in\{1,2,...,j-2\}$, history $\mathbf{h}$ belongs to $H(i)$ if and only if it minimizes $h_{1}+...+h_{i}$ among all vectors in $H(i+1)$. This leads to a unique vector in $H(1)$, which we denote by $\mathbf{h^{*}}\equiv(h_{1}^{*},...,h_{l}^{*})$. By construction, $h_{j+1}^{*}=...=h_{l}^{*}=0$ and $h_{j}^{*}=1$.
Strict Monotone Equilibria:
Suppose for some $\delta$ is close enough to $1$, there exists a strict monotone equilibrium under which $j\geq 2$. The monotonicity of the belief map $\pi$ implies that $\pi(\mathbf{h^{*}})\leq\pi(\mathbf{h})$ for any $\mathbf{h}\in H(j-1)$. We establish the following lemma, which shows that $\mathbf{h^{*}}$ can only originate from $\mathbf{h^{*}}-\mathbf{1}_{j}$.
@@
Fix any strict monotone equilibrium. If an agent’s history is $\mathbf{h^{*}}$, then his immediate predecessor’s history is either $\mathbf{h^{*}}$ or $\mathbf{h^{*}}-\mathbf{1}_{j}$.
The proof is in Appendix A.4.
For every $\mathbf{h}\in H_{\sigma}$, let $D(\mathbf{h})\equiv\{i:\sigma_{D}(\mathbf{h},s_{i})\neq\mathbf{h}\}\subset\{1,2,...,l\}$. Given the no turning-back property, this implies that for any $\mathbf{h}\in H_{\sigma}$, $i\in D(\mathbf{h})$ if and only if the disclosed vector after receiving private signal $s_{i}$ at history $\mathbf{h}$ is not weakly less than $\mathbf{h}$, or equivalently, the number of signal $s_{i}$ disclosed is more than the $i$th entry of $\mathbf{h}$. For any $\mathbf{h}=(h_{1},...,h_{l})$ with $h_{i}\neq 0$ for some $i\geq 2$, we have $D(\mathbf{h})\neq\emptyset$. This is because $\pi$ being monotone implies that $\pi(\mathbf{h}-\mathbf{1}_{i}+\mathbf{1}_{1})>\pi(\mathbf{h})$, so $1\in D(\mathbf{h})\neq\emptyset$. Whenever the set $D(\mathbf{h})$ is non-empty, let $\overline{d}(\mathbf{h})\equiv\max D(\mathbf{h})$.
Since the equilibrium is strict, we know that $\pi(\mathbf{h})>\pi(\mathbf{h^{\prime}})$ for every $\mathbf{h}\in H_{\sigma}$ and $\mathbf{h^{\prime}}<\mathbf{h}$. Hence, if the agent receives private signal $s_{i}$ at history $\mathbf{h}$ with $i\notin D(\mathbf{h})$, he will disclose history $\mathbf{h}$ to his immediate successor, so the history will remain the same conditional on communication succeeds. This together with Lemma 2 leads to